multiply (2x-3)(4x exponent of 2+5x-6)
Answer by midwood_trail(310)

(2x - 3) (4x^2 + 5x - 6)
You are multiplying a binomial by a trinomial.
Multiply one term at a time.
We start by multiplying the trinomial by 2x.
2x times 4x^2 = 8x^3
2x times 5x = 10x^2
2x times - 6 = -12x
We now do the same thing using -3.
-3 times 4x^2 = -12x^2
-3 times 5x = -15x
-3 times - 6 = 18
We now put evertything together and simplify by combining like terms.
8x^3 + 10x^2 - 12x - 12x^2 - 15x + 18
8x^3 -2x^2 - 27x + 18
That's it.
Did you follow?